# Mohammad Mehdi Morovati    Florian Tambon    Mina Taraghi    Amin Nikanjam    Foutse Khomh

### v2

@@ 1 Introduction
Although there exist studies on testing and debugging RL programs zolfagharian2022search ; tambon2023mutation , the main challenges and obstacles that developers face while developing RL applications are still  zhang2020deep . Moreover, because of basic differences between the paradigm of traditional software applications and ML applications morovati2023bugs ; islam2020repairing , it is expected that developers of ML applications face different types of challenges in the implementation process of such applications. Thus, DRL developers may face different challenges from other types of software systems (including traditional software systems as well as other subcategories of ML applications) nguyen2020deep ; du2021survey ; dulac2021challenges . As an example, Listing.1 shows a SO post (70562317) related to a DRL application, representing a challenge in implementing the method to choose an optimal action which is specific to DRL development and differs from ML and DL development challenges.
@@
84SELECT * FROM Posts
85WHERE Tags LIKE ’%<deep-learning>%’
86AND Tags LIKE ’%<reinforcement-learning>%’
@@ 1.1 Manual Inspection
During this step, a team of four raters (three Ph.D. candidates and a senior research staff who all are practitioners of DRL development) is responsible for labeling the collected SO posts. Following a methodology similar to prior works humbatova2020taxonomy ; islam2019comprehensive , we split the collected SO posts into 10 parts each inspected in a dedicated labeling round. All the discussions and referenced source codes in each post were thoroughly reviewed. The raters use an open coding method lune2017qualitative to label SO posts and categorize them. Each SO post is reviewed by two raters. We use the “Google Sheets” platform googleSheet to save all extracted labels in an online environment. That is, all raters put generated labels in the shared document, but they do not have access to the labels that other raters assigned to each post. After labeling SO posts in each round, the raters meet to discuss disagreements and resolve conflicts. In the case they fail to resolve a disagreement, a third rater reviews the SO post and makes a decision about its label, acting as a tie-breaker. Besides, the raters review the generated labels in each meeting to ensure their comprehensiveness and granularity (combining similar labels generated by different raters or dividing a label into separate ones). The justification for dividing SO posts into 10 labeling rounds, with approximately 10% of the total collected posts in each round, emerges from the necessity for raters’ discussion in iterative meetings. These discussions are essential for resolving labeling conflicts and reaching an agreement, reviewing generated labels, and creating finer or coarser-grained labels. In case of changing existing labels, raters re-review the previously labeled posts to ensure that assigned labels are in synchronization with newly generated labels. We made this decision to allow for continuous improvement of the labeling process. This way raters have the opportunity to resolve their conflicts at the end of each round, similar to the technique used in previous studies humbatova2020taxonomy ; islam2019comprehensive . Besides, in the case that any of the raters suggests generating a new label, all raters meet to discuss and reach an agreement on that new label. After completing labeling all 927 posts, all raters meet to finalize the generated labels, categorize them, and create the taxonomy. Then, the first two authors review all of the labeled posts again to ensure that their assigned labels are in sync with the final generated taxonomy. Regarding the posts in which the questioner asks more than one question belonging to multiple categories of challenges, we repeat that record in our dataset and assign a different label to each record. For instance, post #45382763 has been identified to belong to two categories, comprehension and design problems. Although we could not report inter-rater agreement level due to the lack of prior defined categories, we calculate inter-rater agreement between the pair of raters who investigate each SO post after finalizing the labels using Cohen’s Kappa mcdonald2019reliability and obtained an 86% agreement level. Table 1 presents detailed information on the labeling procedure.
@@
It is worth mentioning that 70% of the DRL-related questions in SO still remain without any accepted answer. This is consistent with previous findings by Alshangiti et al. alshangiti2019developing that 61% of ML-related SO posts remain without any accepted answer. Multiple factors could explain this finding. In some cases, the person who asks the question responds to the question after a while, but she does not assign the accepted answer badge to the post (e.g., #45364837). Some users also ask basic questions irrelevant to DRL but assign DRL-related tags to them. These questions receive negative scores and remain without any accepted answer (e.g., #50544568). We also observed posts where the person asking the question forgot to assign the accepted answer badge to an answer, based on upvotes to the response, comments of the person who asked the question, or other people with the same problem under the response (e.g., #63250935). About the posts with accepted answers, it should be mentioned that 16% of them have been answered by the user who published the question. This usually happens when a user asks a very specific question that remains unanswered for a long time, and then the same user finds the response elsewhere and adds it to his original post (e.g., #2723999).
@@ 1.2 Taxonomy Construction and Validation
The comprehensiveness and representativeness of the obtained taxonomy are assessed using a survey with DRL developers/practitioners who are not involved in the construction of the taxonomy. While we built our taxonomy based on SO posts, we used Github to identify potential respondents to our survey. We collect a list of survey participants from collaborators of GitHub repositories related to DRL. Specifically, we extract GitHub repositories mentioning ’deep reinforcement learning’ in their description using GitHub’s search API V3; a rest API that receives a query and returns a list of repositories that satisfy conditions stated in the query. In other words, we use ‘deep reinforcement learning’ as the keyword to generate the search query of GitHub search API V3111https://docs.github.com/en/rest. Given that GitHub search API limits access to only the first 1,000 results, we follow the methodology used in morovati2023bugs and run several different queries to achieve less than 1,000 repositories for each query. That is, we divide the duration of the search for repository creation date between Jan 1, 2010, and Jan 31, 2023 (the date of running queries) into snapshots of 1 month. Thus, we execute 157 GitHub search requests to collect 7,244 repositories. Subsequently, we filter out forked and disabled repositories. The complete list of repositories and a sample search query to extract DRL-related repositories are accessible via the replication package replication_package . In the next step, we check the repositories’ contributors and collect the contributors mentioning their email addresses, obtaining 2531 unique email addresses of developers.
@@
We use Qualtrics survey_ualtrics , an online survey tool for designing and conducting surveys, to create survey forms. Table 2 presents the structure of our survey questionnaires. The survey starts with general questions regarding the participant’s current role and experience in DRL development (Section 1 of Table 2). Subsequently, we delve into specific questions regarding each DRL development challenge which is pointed out in the finalized taxonomy (Section 2 of Table 2). Concerning the potential complexity and difficulty of comprehending the whole taxonomy as a single figure within the survey, we present challenges (subcategories) in groups based on their respective main categories. Besides, to ensure clarity, a detailed description has been given with each challenge offering participants a thorough understanding of the asked challenges. For each identified challenge, we ask three questions including 1) a ‘yes/no’ question identifying whether the answerer has faced the identified challenge, 2) the severity of the challenge, and 3) the amount of effort required to address the challenge. If a participant answers ‘yes’ to the first question, s/he will have the next two Likert-scale questions which are related to the severity of challenges and the required effort to address them. We also provide a free-text question in the final part of the survey asking the participants about any challenges in developing DRL applications not listed in our provided taxonomy (Section 3 of Table  2). These free-text questions allow us to collect possible challenges that we may have missed in the taxonomy. The full survey questionnaire is available in our replication package replication_package .
@@ 2 Results
DRL Issues. This category focuses on the challenges that developers may face while developing the DRL part of their applications. Challenges that belong to this category are specific to DRL application development. That is, compared to challenges classified under other categories that can be shared among other kinds of ML-related applications (e.g. supervised learning), challenges of the DRL issues category only are faced during DRL application development.
@@
    Design Problem: Instances where the user asks for advice for designing a solution and implementing a DRL application for their specific problem or scenario. For example, developers asked for recommendations to implement different parts of DRL applications for Curve Fever or mini-golf games. Another such challenge in this category is related to designing the properties of each object in a tank game.
@@
      Train: This subcategory includes questions regarding the theoretical concepts of the DL model part and its training in DRL applications. Training part is responsible for training the agent in order to select optimal actions in various situations.
@@
      Problem attributes: SO posts posing theoretical queries about attributes of DRL problems (such as reward, action, state, etc.) are categorized in this subcategory.
@@
      Algorithm: Questions on understanding basic concepts of various DRL algorithms, e.g., actor-critic, q-learning, SARSA, etc. are classified as algorithm subcategories. To develop a clear representation of the most questioned DRL algorithms, we categorized this subcategory into different DRL algorithms.
@@
    Reward: Challenges in the implementation of reward, e.g., not using the negative reward to penalize each added time step. Another such example of this category is a post asking “I am implementing the basic RL algorithm to play the game Flappy Bird. I want to be able to process the screen and recognize whether a point has been scored or the bird has died. Processing the screen returns a stacked numpy array. The reward function then needs to assign a reward to the provided array, but I have no idea how to go about this”.
@@
    State/Observation: Questions/Problems regarding the state(s) or the agent observation(s), e.g., handling large state spaces. An example of such instance is a user asking “I implemented a 3x3 OX game by q-learning (it works perfectly in AI v.s AI and AI v.s Human), but I can’t go one step further to 4x4 OX game since it will eat up all my PC memory and crash…Since I need to calculate each Q value (for each state, each action), I need such a large number of array, is it expected? any way to avoid it?” and the accepted answer provided them with suggestions for reducing their state space size by considering symmetries and other tricks.
@@
    Hyperparameters: Questions/Problems related to the hyperparameters of the RL algorithm, e.g., setting the discount factor too high. A good example of this group is demonstrated in a post where the user shared the code for their learning algorithm and reported that the loss keeps increasing and the model is not learning. The accepted answer states that “The main problem I think is the discount factor, gamma. You are setting it to 1.0, which means that you are giving the same weight to the future rewards as the current one”.
@@
    Dependency: This subcategory includes questions/problems about the mismatch between versions of installed libraries/frameworks and problems in installed versions of libraries, e.g., when the version of the installed OpenAI Gym is not compatible with Python. An instance of this is a user reporting getting an error installing OpenAIGym and the answer pointed out that “the error means that the package has dependency requirements that conflict with one another”.
@@
    API usage: This subcategory includes questions about the usage of arguments, attributes, methods, etc. of an API. It also includes questions about the default values, implemented method, existence of attributes, or methods in an API. An example from this group of issues is a user reporting not knowing how to get the weights of the network using the correct API methods: “I’m using RLlib to train a reinforcement learning policy (PPO algorithm). I want to see the weights in the neural network underlying the policy. After digging through RLlib’s PPO object, I found the TensorFlow Graph object. I thought that I would find the weights of the neural network there. But I can’t find them”.
@@
      API misuse: This subcategory covers SO posts that mentioned misunderstanding API usage. In other words, API misuse occurs when DRL developers try to utilize an API in a manner that is not aligned with its intended purpose.
@@
  3. 2.
    Data Preprocessing: Questions about preparing data to be fed into the DL model, e.g., the shape of the input matrix. As an example there was a user that asked “I am learning how to use Gym environments to train deep learning models built with TFLearn. At the moment my array of observations has the following shape: (210, 160, 3). Any recommendations on what is the best way to reshape this array so it can be used in a TensorFlow classification model?”
  4. 3.
    DL framework: Questions about the usage of DL frameworks (e.g., Keras, TensorFlow, PyTorch, etc.) in development of DL part in DRL applications. For instance, a user who wanted to use Huber loss in their model which was written using Keras, but this loss function is not readily available in this library, and the accepted answer implements the Huber loss function.
@@
  3. 2.
    Distributed processing: Questions/Problems about running DRL applications as a distributed software. For example when a user wanted to implement the Asynchronous Advantage Actor Critic (A3C) model for reinforcement learning in their local machine, they posed a question about the possibility of implementing it in a distributed manner: “Would it be easier/faster/better to implement this using the distributed TensorFlow API? In the documentation and talks, they always make explicit mention of using it in multi-device environments. I don’t know if it’s an overkill to use it in a local async algorithm”.
  4. 3.
    Multi-threading: Questions/Problems about running DRL applications as a multi-threaded software. An example of this category was observed in a post where the user asked about the possibility of reducing the training time by running it on multiple threads concurrently: “My friend and I are training a DDQN for learning 2D soccer. I trained the model about 40.000 episodes but it tooks 6 days. Is there a way for training this model concurrently? For example, I have 4 core and 4 thread and each thread trains the model 10.000 times concurrently. Therefore, time to training 40.000 episodes are reduced 6 days to 1,5 days like parallelism of for loop”.
  5. 4.
@@
Fig.2 presents the distribution of SO posts related to DRL application development for a period of 13 years, from 2009 to 2022. From Fig.1(a), we observe a substantial surge of inquiries about DRL development in 2016; reaching a peak in 2019 and 2020. Additionally, as can be seen on Fig.1(b), comprehension and design problem questions dominated posts about DRL application development challenges. It is also noticeable that API usage, the second most common DRL application development challenge, was at its peak in 2018.
@@
Fig.3 presents the distribution of time taken by SO posts from different categories to receive an accepted answer. This duration is an indicator of the level of difficulty of the questions contained in the SO posts in the development of both traditional  haque2020challenges ; zahedi2020mining and ML software alshangiti2019developing ; chen2020comprehensive . Parallel Processing is the category with the highest average time taken before receiving an accepted answer. This can be explained by the fact that using multi-processing or distributed processing in DRL is not necessarily widespread and also requires particular knowledge and expertise. The remaining categories need a nearly similar average time frame to receive an accepted answer, with the general programming issues category having the shortest average duration. We attribute the shortest average time of the general programming issues category to the fact that this category contains generic challenges that do not require expertise in DL or DRL.
@@
Although we note a high proportion of outliers in Fig.3, we have a relatively low median, first, and third quartiles, for all categories compared to the average. Meaning that, although the majority of posts received an accepted answer in a relatively short period of time (in general less than 10 days), a sizable number of posts required a longer time (more than 20 days). Therefore, even within categories, there is quite some discrepancy inside the subcategories (challenges) themselves. As can be seen on Fig.3(a), SO posts categorized as Model have a higher average waiting time to receive an accepted answer than the other subcategories of the DL Issues category. Among subcategories in DRL Issues, hyperparameters includes SO posts with the highest average required time for receiving an accepted answer (Fig.3(b)). For DRL libraries/frameworks, SO posts belonging to API usage, Dependency and Installation subcategories require the longest average time before receiving an accepted answer (Fig.3(c)). Regarding subcategories within parallel Processing & multi-threading, it is notable that the average duration required to receive an accepted answer for SO posts classified under multi-processing subcategory exceeds one year. It should be also taken into account that the small number of SO posts in this subcategory might bias the results, with respect to the fact that small data may not represent the distribution of classes in the population adequately bruer2015designing .
@@
We cross-checked the taxonomy generated based on SO posts using a validation survey. 65 ML practitioners participated in our survey to assess our identified challenges, including 55% researchers (Master’s and Ph.D. students, research assistants, and professors), 27% ML/SE engineers, 11% developers, and 7% data scientists. Among our respondents, 86% have at least 1 year of DRL development experience and 46% have more than 3 years of experience. Table 4 summarizes the responses of participants for each DRL development challenge contained in the taxonomy. For each challenge, we provide the percentage of developers who reported having experienced that challenge (based on the answers to the ‘yes’ or ‘no’ question). Besides, we asked the participants about the severity of each challenge and the effort required to address them. Results show that all challenges presented in our taxonomy were encountered by the survey respondents. Moreover, no additional challenges were proposed by the survey participants through the open-text questions. This indicates that our provided taxonomy is representative of challenges faced by developers during the DRL application development.

### v3

@@ 1 Introduction
Although there exist studies on testing and debugging RL programs zolfagharian2022search ; tambon2023mutation , the main challenges and obstacles that developers face while developing RL applications are still unclear zhang2020deep . Moreover, because of basic differences between the paradigm of traditional software applications and ML applications morovati2023bugs ; islam2020repairing , it is expected that developers of ML applications face different types of challenges in the implementation process of such applications. Thus, DRL developers may face different challenges from other types of software systems (including traditional software systems as well as other subcategories of ML applications) nguyen2020deep ; du2021survey ; dulac2021challenges .
As an example, Listing.1 shows a SO post (70562317) related to a DRL application, representing a challenge in implementing the method to choose an optimal action which is specific to DRL development and differs from ML and DL development challenges.
@@
85SELECT * FROM Posts
86WHERE Tags LIKE ’%<deep-learning>%’
87AND Tags LIKE ’%<reinforcement-learning>%’
@@ 1.1 Manual Inspection
During this step, a team of four raters (three Ph.D. candidates and a senior research staff who all are practitioners of DRL development) is responsible for labeling the collected SO posts. Following a methodology similar to prior works humbatova2020taxonomy ; islam2019comprehensive , we split the collected SO posts into 10 parts, each of which is inspected in a dedicated labeling round. All the discussions and referenced source codes in each post are thoroughly reviewed. The raters use an open coding method lune2017qualitative to label SO posts and categorize them. Each SO post is reviewed by two raters. We use the “Google Sheets” platform googleSheet to save all extracted labels in an online environment. That is, all raters put generated labels in the shared document, but they do not have access to the labels that other raters assigned to each post. After labeling SO posts in each round, the raters meet to discuss disagreements and resolve conflicts. In the case they fail to resolve a disagreement, a third rater reviews the SO post and makes a decision about its label, acting as a tie-breaker. Besides, the raters review the generated labels in each meeting to ensure their comprehensiveness and granularity (combining similar labels generated by different raters or dividing a label into separate ones). The justification for dividing SO posts into 10 labeling rounds, with approximately 10% of the total collected posts in each round, emerges from the necessity for raters’ discussion in iterative meetings. These discussions are essential for resolving labeling conflicts and reaching an agreement, reviewing generated labels, and creating finer or coarser-grained labels. In case of changing existing labels, raters re-review the previously labeled posts to ensure that assigned labels are in synchronization with newly generated labels. We made this decision to allow for continuous improvement of the labeling process. This way raters have the opportunity to resolve their conflicts at the end of each round, similar to the technique used in previous studies humbatova2020taxonomy ; islam2019comprehensive . Besides, in the case that any of the raters suggests generating a new label, all raters meet to discuss and reach an agreement on that new label. After completing labeling all 927 posts, all raters meet to finalize the generated labels, categorize them, and create the taxonomy. Then, the first two authors review all of the labeled posts again to ensure that their assigned labels are in sync with the final generated taxonomy. Regarding the posts in which the questioner asks more than one question belonging to multiple categories of challenges, we repeat that record in our dataset and assign a different label to each record. For instance, post #45382763 has been identified to belong to two categories, comprehension and design problems. Although we could not report inter-rater agreement level due to the lack of prior defined categories, we calculate inter-rater agreement between the pair of raters who investigate each SO post after finalizing the labels using Cohen’s Kappa mcdonald2019reliability and obtained an 86% agreement level. Table 1 presents detailed information on the labeling procedure.
@@
It is worth mentioning that 70% of the DRL-related questions in SO still remain without any accepted answer. This is consistent with previous findings by Alshangiti et al. alshangiti2019developing that 61% of ML-related SO posts remain without any accepted answer. Multiple factors could explain this finding. In some cases, the person who asks the question responds to the question after a while, but she does not assign the accepted answer badge to the post (e.g., #45364837). Some users also ask basic questions irrelevant to DRL but assign DRL-related tags to them. These questions receive negative scores and remain without any accepted answer (e.g., #50544568). We also observe posts where the person asking the question forgot to assign the accepted answer badge to an answer, based on upvotes to the response, comments of the person who asked the question, or other people with the same problem under the response (e.g., #63250935). About the posts with accepted answers, it should be mentioned that 16% of them have been answered by the user who published the question. This usually happens when a user asks a very specific question that remains unanswered for a long time, and then the same user finds the response elsewhere and adds it to his original post (e.g., #2723999).
@@ 1.2 Taxonomy Construction and Validation
Interviews and surveys stand as the two popular methods to validate the results of qualitative studies hove2005experiences ; aldhaen2020interview . Considering the advantages of conducting surveys, including cost-effectiveness, generalizability, reliability, and versatility decarlo2018scientific ; nekkanti2016surveys , we assessed the comprehensiveness and representativeness of the obtained taxonomy using a survey with DRL developers/practitioners who are not involved in the construction of the taxonomy. Nevertheless, it is noteworthy that several preceding studies have presented their findings without undergoing any validation process zhang2019empirical ; islam2019comprehensive .
@@
While we build our taxonomy based on SO posts, we use Github to identify potential respondents to our survey. We collect a list of survey participants from collaborators of GitHub repositories related to DRL. Specifically, we extract GitHub repositories mentioning ’deep reinforcement learning’ in their description using GitHub’s search API V3; a rest API that receives a query and returns a list of repositories that satisfy conditions stated in the query. In other words, we use ‘deep reinforcement learning’ as the keyword to generate the search query of GitHub search API V3111https://docs.github.com/en/rest. Given that GitHub search API limits access to only the first 1,000 results, we follow the methodology used in morovati2023bugs and run several different queries to achieve less than 1,000 repositories for each query. That is, we divide the duration of the search for repository creation date between Jan 1, 2010, and Jan 31, 2023 (the date of running queries) into snapshots of 1 month. Thus, we execute 157 GitHub search requests to collect 7,244 repositories. Subsequently, we filter out forked and disabled repositories. The complete list of repositories and a sample search query to extract DRL-related repositories are accessible via the replication package replication_package . In the next step, we check the repositories’ contributors and collect the contributors mentioning their email addresses, obtaining 2531 unique email addresses of developers.
@@
We use Qualtrics survey_ualtrics , an online survey tool for designing and conducting surveys, to create survey forms. Table 2 presents the structure of our survey questionnaires. The survey starts with general questions regarding the participant’s current role and experience in DRL development (Section 1 of Table 2). Subsequently, we delve into specific questions regarding each DRL development challenge which is pointed out in the finalized taxonomy (Section 2 of Table 2). Concerning the potential complexity and difficulty of comprehending the whole taxonomy as a single figure within the survey, we present challenges (subcategories) in groups based on their respective main categories. Besides, to ensure clarity, a detailed description is given with each challenge offering participants a thorough understanding of the asked challenges. For each identified challenge, we ask three questions including 1) a ‘yes/no’ question identifying whether the answerer has faced the identified challenge, 2) the severity of the challenge, and 3) the amount of effort required to address the challenge. If a participant answers ‘yes’ to the first question, s/he will have the next two Likert-scale questions which are related to the severity of challenges and the required effort to address them. We also provide a free-text question in the final part of the survey asking the participants about any challenges in developing DRL applications not listed in our provided taxonomy (Section 3 of Table  2). These free-text questions allow us to collect possible challenges that we may miss in the taxonomy. The full survey questionnaire is available in our replication package replication_package .
We conduct a comprehensive analysis of top DRL-related GitHub repositories to ensure the completeness of the generated taxonomy. Besides, investigating the challenges faced by developers engaged in the development of real-world DRL applications may further enhance the generalizability of the generated taxonomy. Our methodology for selecting DRL-related repositories aligns with established approaches documented in similar prior studies morovati2023bugs ; morovati2024bug ; humbatova2020taxonomy . Initially, we extracted 7,244 repositories to identify contributors of DRL-related repositories. Then, we select the top 100 repositories based on the highest number of stars. Subsequently, we conduct searches for ‘challenge’, ‘difficult’, and ‘complex’ keywords within repositories’ documentation, commit messages, and closed issues to identify potential challenges encountered by developers during the development of these repositories. Our search yielded 746 occurrences of the aforementioned keywords within the top 100 DRL-related repositories. Using a confidence level of 95% and a confidence interval of 5%, we conduct sampling, resulting in the selection of 254 instances. In the next step, two raters independently examine 254 randomly selected commits and issues to find out whether they pertained to real DRL-related development challenges. The raters carefully review commit messages, issue discussions, and investigate any changes made in relation to the specified keywords.
@@ 2 Results
DRL Issues. This category focuses on the challenges that developers may face while developing the DRL part of their applications. Challenges that belong to this category are specific to DRL application development. That is, compared to challenges classified under other categories that can be shared among other kinds of ML-related applications (e.g. supervised learning), challenges of the DRL issues category are only faced during DRL application development.
@@
    Design Problem: Instances where the user asks advice for designing a solution and implementing a DRL application for their specific problem or scenario. For example, developers asked for recommendations to implement different parts of DRL applications for Curve Fever or mini-golf games. Another such challenge in this category is related to designing the properties of each object in a tank game. Although we tried to make finer-grained subcategories under this subcategory, we found out that it is impractical to split this subcategory further, as the majority of this subcategory’s posts primarily comprised users seeking guidance on high-level conceptual queries about defining their problem within a DRL context.
@@
      Training: This subcategory comprises inquiries concerning the theoretical concepts of the DL model and its significance in DRL applications. The DL model plays an essential role in the training and decision-making process of the DRL agents. Indeed, DL models in DRL applications serve to represent policy of the agent, estimate the value of state-action, and learn mappings from states to values.
@@
      Problem attributes: SO posts containing theoretical queries about attributes of DRL problems (such as reward, action, state, etc.) are categorized in this subcategory. SO posts in this category are related to the conceptual aspects of reward, action, and state, not their implementation details. For example, questions on how to formulate states, actions, and the reward signal for a particular DRL problem, or why a particular definition of the states is not suitable for a problem.
@@
      Algorithm: Questions related to fundamental concepts of various DRL algorithms, such as actor-critic, Q-learning, SARSA, etc., are categorized as algorithm subcategories. The main duty of DRL algorithms is updating parameters of the policy and value function according to the observed states and achieved rewards. Our investigation into these algorithms reveals that Q-learning (28%) and DQN (15.3%) are the most common DRL-related algorithms posing challenges for developers. Conversely, Actor-Critic (3.1%), Proximal Policy Optimization (2.6%), and Trust Region Policy Optimization (0.5%) are the least queried DRL-related algorithms.
@@
    Reward: Challenges in the implementation of reward, e.g., not using the negative reward to penalize each added time step are categorized in this subcategory. An example of this category is a post asking “I am implementing the basic RL algorithm to play the game Flappy Bird. I want to be able to process the screen and recognize whether a point has been scored or the bird has died. Processing the screen returns a stacked numpy array. The reward function then needs to assign a reward to the provided array, but I have no idea how to go about this”.
@@
    State/Observation: Questions/Problems regarding the state(s) or the agent observation(s), e.g., handling large state spaces. An example of such instance is a user asking “I implemented a 3x3 OX game by q-learning (it works perfectly in AI v.s AI and AI v.s Human), but I can’t go one step further to 4x4 OX game since it will eat up all my PC memory and crash…Since I need to calculate each Q value (for each state, each action), I need such a large number of array, is it expected? any way to avoid it?” and the accepted answer offered suggestions on reducing their state space size by considering symmetries and other tricks.
@@
    Hyperparameters: Questions/Challenges related to the hyperparameters of the RL algorithm, e.g., setting the discount factor too high. A good example of this group is demonstrated in a post where the user shared the code for the learning algorithm and reported that the loss keeps increasing and the model is not learning. The accepted answer states that “The main problem I think is the discount factor, gamma. You are setting it to 1.0, which means that you are giving the same weight to the future rewards as the current one”.
@@
    Dependency: This subcategory includes questions/challenges about the mismatch between versions of installed libraries/frameworks and problems in installed versions of libraries, e.g., when the version of the installed OpenAI Gym is not compatible with Python. An instance of this subcategory is a user reporting getting an error while installing OpenAIGym and the answer pointed out that “the error means that the package has dependency requirements that conflict with one another”.
@@
    API usage: This subcategory includes questions about the usage of arguments, attributes, methods, etc. of an API. It also includes questions about the default values, implemented method, existence of attributes, or methods in an API. An example from this group of issues is a user reporting not knowing how to get the weights of the network using the correct API methods: “I’m using RLlib to train a reinforcement learning policy (PPO algorithm). I want to see the weights in the neural network underlying the policy. After digging through RLlib’s PPO object, I found the TensorFlow Graph object. I thought that I would find the weights of the neural network there. But I can’t find them”. This subcategory is subdivided into five subcategories to delineate more detailed challenges.
@@
      API misuse: This subcategory covers SO posts that mention misunderstanding of API usage. In other words, API misuse occurs when DRL developers try to utilize an API in a manner that is not aligned with its intended purpose.
@@
  2. b.
    Data Preprocessing: Questions about preparing data to be fed into the DL model, e.g., the shape of the input matrix. As an example, there was a user who asked “I am learning how to use Gym environments to train deep learning models built with TFLearn. At the moment my array of observations has the following shape: (210, 160, 3). Any recommendations on what is the best way to reshape this array so it can be used in a TensorFlow classification model?”
  3. c.
    DL framework: Questions about the usage of DL frameworks (e.g., Keras, TensorFlow, PyTorch, etc.) in development of DL part in DRL applications. For instance, a user who wanted to use Huber loss in their model which was written using Keras, but this loss function is not readily available in Keras, and the accepted answer implements the Huber loss function.
@@
  2. b.
    Distributed processing: Questions/Challenges about running DRL applications as a distributed software. For example when a user wanted to implement the Asynchronous Advantage Actor Critic (A3C) model for reinforcement learning in the local machine, she posed a question about the possibility of implementing it in a distributed manner: “Would it be easier/faster/better to implement this using the distributed TensorFlow API? In the documentation and talks, they always make explicit mention of using it in multi-device environments. I don’t know if it’s an overkill to use it in a local async algorithm”.
  3. c.
    Multi-threading: Questions/Problems about running DRL applications as a multi-threaded software. An example of this category was observed in a post where the user asked about the possibility of reducing the DRL application’s training time by running it on multiple threads concurrently: “My friend and I are training a DDQN for learning 2D soccer. I trained the model about 40.000 episodes but it tooks 6 days. Is there a way for training this model concurrently? For example, I have 4 core and 4 thread and each thread trains the model 10.000 times concurrently. Therefore, time to training 40.000 episodes are reduced 6 days to 1,5 days like parallelism of for loop”.
  4. d.
@@
Table 3 shows the number and percentage of the SO posts within our dataset having different tags and/or keywords that we used in subsection LABEL:subsec:so_extract. Fig.2 presents the distribution of SO posts related to DRL application development for a period of 13 years, from 2009 to 2022. From Fig.2(a), we observe a substantial surge of inquiries about DRL development in 2016; reaching a peak in 2019 and 2020. Additionally, as can be seen on Fig.2(b), comprehension and design problem questions dominated posts about DRL application development challenges. It is also noticeable that API usage, the second most common DRL application development challenge, was at its peak in 2018.
@@
Fig.3 depicts the distribution of time taken by SO posts from different categories to receive an accepted answer. This duration is an indicator of the difficulty level of the questions mentioned in the SO posts in the development of both traditional  haque2020challenges ; zahedi2020mining and ML software alshangiti2019developing ; chen2020comprehensive . Parallel Processing is the category with the highest average time taken before receiving an accepted answer. This can be explained by the fact that using multi-processing or distributed processing in DRL is not necessarily widespread and also requires particular knowledge and expertise. The remaining categories need a nearly similar average time frame to receive an accepted answer, with the general programming issues category having the shortest average duration. We attribute the shortest average time of the general programming issues category to the fact that this category contains generic challenges that do not require expertise in DL or DRL.
@@
Although we note a high proportion of outliers in Fig.3, we have a relatively low median, first, and third quartiles, for all categories compared to the average. Meaning that, although the majority of posts received an accepted answer in a relatively short period of time (in general less than 10 days), a sizable number of posts required a longer time (more than 20 days). Therefore, even within categories, there is quite some discrepancies inside the subcategories (challenges) themselves. As can be seen on Fig.4(a), SO posts categorized as Model have a higher average waiting time to receive an accepted answer than the other subcategories of the DL Issues category. Among subcategories under DRL Issues, hyperparameters includes SO posts with the highest average required time for receiving an accepted answer (Fig.4(b)). About DRL libraries/frameworks, SO posts belonging to API usage, Dependency and Installation subcategories require the longest average time before receiving an accepted answer (Fig.4(c)). Regarding subcategories within parallel Processing & multi-threading, it is notable that the average duration required to receive an accepted answer for SO posts classified under multi-processing subcategory exceeds one year. It should be also taken into account that the small number of SO posts in this subcategory might bias the results, with respect to the fact that small data may not represent the distribution of classes in the population adequately bruer2015designing .
Although our methodology to measure the difficulty level of addressing challenges aligns with prior studies analyzing SO posts alshangiti2019developing ; decan2019empirical , it is worth noting that some studies have utilized the number of posts within each category as a metric to illustrate the difficulty associated with addressing the related challenges bangash2019developers ; hamidi2021towards . In the case of considering the number of posts as an indicator of the difficulty level for tackling challenges (the number mentioned along with each category in Fig. 1), the results would closely mirror the difficulty levels observed for the challenges categorized as DL issues and parallel processing categories depicted in Fig.4. For instance, design problem is the second most difficult challenge in the DRL issues category when we use ‘number of posts’ or ‘duration to receive an accepted answer’ to measure the difficulty level. Conversely, the difficulty levels for challenges falling under DRL issues and DRL libraries/frameworks differ by using these two metrics. For example, as illustrated in Fig.4(b), hyperparameters emerges as the most challenging subcategory within the DRL issues category, even though comprehension has the highest number of posts. This discrepancy may be attributed to the nature of inquiring about hyperparameters, which may necessitate various implementations and a longer time to respond compared to other challenges. Furthermore, addressing SO posts categorized as comprehension mainly seeks users’ background knowledge, which does not necessarily require practical implementation or application execution.
@@
We cross-check the taxonomy generated based on SO posts using a validation survey. 65 ML practitioners participated in our survey to assess our identified challenges, including 55% researchers (Master’s and Ph.D. students, research assistants, and professors), 27% ML/SE engineers, 11% developers, and 7% data scientists. Among our respondents, 86% have at least 1 year of DRL development experience and 46% have more than 3 years of experience. Table 4 summarizes the responses of participants for each DRL development challenge contained in the taxonomy. For each challenge, we provide the percentage of developers who reported having experienced that challenge (based on the answers to the ‘yes’ or ‘no’ question). Besides, we asked the participants about the severity of each challenge and the effort required to address them. Results show that all challenges presented in our taxonomy were encountered by the survey respondents. Moreover, no additional challenges were proposed by the survey participants through the open-text questions. This indicates that our provided taxonomy is representative of challenges faced by developers during the DRL application development. It is worth noting that the survey questions incorporate the challenges depicted in the third level of the taxonomy (Fig.1), ensuring the survey’s conciseness.